# Learning to Rank Critical Road Segments via Heterogeneous Graphs with OD Flow Integration

### v2

@@ Abstract
Existing learning-to-rank methods for road networks often fail to incorporate origin–destination (OD) flows and route information, limiting their ability to model long-range spatial dependencies. To address this gap, we propose HetGL2R, a heterogeneous graph learning framework for ranking road-segment importance. HetGL2R builds a tripartite graph that unifies OD flows, routes, and network topology, and further introduces attribute-guided graphs that elevate node attributes into explicit nodes to model functional similarity. A heterogeneous joint random walk algorithm (HetGWalk) samples both graph types to generate context-rich node sequences. These sequences are encoded with a Transformer to learn embeddings that capture long-range structural dependencies driven by OD demand and route configuration, as well as functional associations derived from attribute similarity. Finally, a listwise ranking strategy with a KL-divergence loss evaluates and ranks segment importance. Experiments on three SUMO-generated simulated networks of different scales show that, against state-of-the-art methods, HetGL2R achieves average improvements of approximately 7.52%, 4.40% and 3.57% in ranking performance.
@@ 1 Introduction
Efficient and resilient road networks are essential for ensuring smooth urban mobility and public safety. When a single road segment becomes congested or blocked, the resulting disruption often propagates along multiple routes, leading to large-scale delays or even citywide paralysis. Therefore, identifying critical road segments—those whose failure would significantly degrade overall network performance—is of great importance for traffic management and infrastructure planning (rf1).
Figure 1: A small simulated road network. The network contains 11 nodes and two OD pairs: $v_{1}$, $v_{4}>$ with 300 vehicles on one path, and $v_{8}$, $v_{7}>$ with 500 vehicles on two paths.
@@
Prior studies on road-network criticality were primarily based on complex network theory, in which centrality measures (e.g., early betweenness centrality (betweenness) and PageRank (rf3), and more recent methods such as IDME (IDME), EPC (EPC), OVED-Rank (rf44), LSS (rf42), and GLC (rf41)) were used to quantify the structural importance of nodes. These approaches are intuitive and easy to interpret but fail to incorporate the rich attribute features and dynamic traffic behaviors associated with each road segment. In reality, a segment’s criticality depends on multiple factors such as traffic volume, number of lanes, and functional hierarchy, all of which are neglected in purely topological metrics. Some studies have attempted to combine centrality measures with attribute features (e.g., study (rf3)), but most rely on manually tuned parameters, limiting their scalability and generalization to large real-world networks.
@@
In recent years, graph neural network (GNN)–based learning methods (rf4; rf8) have attracted increasing attention in the field of node importance assessment. These models integrate node attributes with local topological information through message-passing mechanisms to learn effective node embeddings. However, due to the limited receptive field of GNNs, they struggle to capture the long-range path dependencies commonly present in traffic flows, which limits their performance in critical road segment ranking tasks.
@@
In transportation networks, the global influence of a road segment is predominantly governed by origin–destination (OD) flows, which quantify the travel demand between each OD pair. The OD flows determine which paths are utilized and how traffic load is distributed, thereby shaping each segment’s functional role and criticality (od1; od2; od3). For instance, if a segment lies on the shortest paths of multiple high-demand OD pairs, any incident or restriction there can induce congestion that propagates upstream along the affected OD paths, leading to severe network-wide delays. Conversely, a peripheral segment carrying a similar traffic volume but serving only a few OD flows would have a far more limited systemic impact if disrupted.
@@
Existing graph learning approaches face two fundamental limitations in this task. First, they primarily learn structural representations from the physical topology of road networks while overlooking the functional structures shaped by OD flows. Second, their local neighborhood aggregation mechanism is constrained by network depth, resulting in limited receptive fields. As a result of these limitations, they struggle to capture long-range spatial dependencies and consequently fail to accurately model the impact patterns of local disruptions (e.g., an incident at node $v_{11}$)—where the scope of impact is dictated by the functional roles of segments within OD flows and paths, rather than by physical proximity. Although CRRank (rf1) incorporates OD information to evaluate the functional importance of road segments, its reliance on manually designed aggregation rules and similarity metrics constrains its expressiveness and effectiveness, as it cannot learn complex interaction patterns in an end-to-end manner.
@@
    Methodological Innovation: HetGL2R constructs a tripartite graph (connecting OD pairs, paths, and road segments) to integrate the functional structure of the road network with its physical structure. Simultaneously, it introduces three attribute-node bipartite graphs—one each for OD pairs, paths, and segments—to explicitly model the attribute-driven functional similarities. A joint random walk mechanism operates on these graphs collectively to sample context-rich node sequences, which are then processed by a Transformer encoder to learn the long-range dependencies and functional roles of segments for accurate criticality ranking.
@@
    Theoretical Insight: Our core insight is that spatial dependencies in traffic networks are governed by functional structure rather than physical topology. To capture this structure, we introduce a joint random walk that integrates two complementary perspectives: functional role relationships (OD-path-segment) and attribute-driven functional similarities. Consequently, the model learns the underlying mechanisms of spatial dependency, thereby enabling accurate assessment of nodes’ spatial influence.
@@ 2.1 Evaluation of node importance in networks
Existing research on identifying critical nodes in networks can be broadly categorized into two approaches: centrality-based methods and machine learning-based methods. The limitation of centrality-based methods lies in their inability to fully leverage the multidimensional features of nodes. Although study (rf3) attempts to improve critical node identification in road networks by integrating node centrality with attribute features, its performance heavily depends on manually set topology-feature trade-off parameters. This leads to a lack of generalizability in quantifying the contributions of topology and features, limiting its applicability in large-scale real-world road networks. To address these limitations, LSS (rf42) estimates node importance based on local information, eliminating the need for global knowledge and complex parameter tuning. Study (rf41) proposed a dual-perspective approach that identifies influential nodes by considering inter-cluster connectivity and local neighborhood importance. However, similar to traditional centrality-based methods, these approaches still fail to incorporate node attribute features or dynamic network behaviors, limiting their applicability in more complex real-world scenarios such as road networks. In recent years, researchers have gradually shifted towards machine learning-based methods, which can be divided into two main technical approaches: reinforcement learning frameworks (e.g., FINDER proposed in study (rf4), which utilizes GraphSAGE (rf5) to learn node importance; study (rf6), which combines attribute-aware graph attention networks with double DQN to optimize critical node identification in sparse graphs) and graph representation learning (e.g., TraNode2vec from study (rf7), which assesses node importance by integrating topological embedding clustering and traffic flow; study (rf8), which introduces the MGL2Rank framework that employs multi-graph fusion techniques to jointly model topology and road segment attributes). However, when applying these methods to road network node importance evaluation, they fail to incorporate OD flow and path selection. As we analyzed in the introduction, these two types of information are crucial for assessing node importance in road networks. Therefore, there remains significant room for improvement in the performance of these methods.
@@ 2.2 Learning to rank
Learning to rank (LTR) originated in the field of information retrieval and is a technique for ordering documents, items, or other entities based on their relevance. In recent years, it has been widely applied in various domains, including recommender systems (rf9), sentiment analysis (rf10), and object detection (rf11). Depending on the characteristics of the input and output spaces, LTR methods can be broadly categorized into three types: pointwise, pairwise, and listwise approaches. The pointwise approach (rf12) treats each item independently, formulating the ranking task as a classification or regression problem. However, this method neglects the relative relationships among items. In contrast, the pairwise approach (rf13) transforms the ranking task into a binary classification problem by comparing item pairs to determine their relative order. Prior research (rf14) has introduced graph neural networks into pairwise ranking to approximate the importance of nodes within a network. Although the pairwise method offers improved accuracy over the pointwise method, it still fails to fully capture the global context and relative importance of items within the entire set, limiting its overall performance. The listwise approach (rf15) treats ranking as a holistic optimization problem and has become a focal point of research due to its modeling and computational complexity. For example, one study (rf16) analyzes how list length and loss function selection affect algorithm stability through generalization error bounds, while another (rf17) creatively reformulates ranking as a multi-task regression problem, overcoming the limitations of traditional linear models. With the rise of deep learning, researchers (rf18; rf19; rf20) have incorporated attention mechanisms and recurrent neural networks into listwise modeling, significantly enhancing performance in scenarios involving complex interactions. On the application side, DiAL (rf21) is the first to integrate the $\alpha$-NDCG diversity metric into listwise loss functions. Further advancements include leveraging large language models to generate reference answer sequences for listwise evaluation (rf22; rf45), as well as developing differentiable approximations to ranking indicators that closely match true loss functions with limited data (rf23). Recent work (rf24) also proposes a regression alignment strategy to address the multi-objective trade-off between score calibration and ranking optimization. Building upon the core principles of listwise ranking, this paper proposes a heterogeneous graph learning framework tailored for node ranking in transportation networks.
@@ 2.3 Heterogeneous graph learning
Heterogeneous graphs (rf25) model multiple types of nodes and edges, integrating information from heterogeneous data sources, thereby demonstrating a stronger representation ability for complex systems. Representation learning on heterogeneous graphs aims to obtain meaningful node embeddings to support various downstream tasks, such as node ranking and link prediction (rf26; rf27). Many studies have focused on designing graph neural network models suitable for heterogeneous graphs. For example, most of these networks (rf28; rf29; rf30) utilize meta-paths and type-aware aggregation strategies to capture the heterogeneity of the graph effectively, that is, the difference in the type of nodes and edges. HetGNN (rf31) combines random walk strategies with bidirectional LSTM structures, effectively integrating heterogeneous node attributes and their neighboring information. HetSANN (rf32) directly encodes the information of different types of nodes and edges through a type-aware attention mechanism, enhancing the model’s ability to represent the heterogeneous graph structure. These methods are distinctive in their modeling strategies, but most still rely on message-passing mechanisms, which do not fully exploit the potential structural relationships between nodes in heterogeneous graphs. To address this, M2CHGNN (rf43) employs a dual-view encoding strategy that integrates attention mechanisms with contrastive learning to generate high-quality node representations for heterogeneous graphs. UI-HEHo (rf47) effectively learns high-order node embeddings via multi-head attention mechanisms and multi-task shared features. SchemaWalk (rf34) proposes an unbiased edge-type exploration strategy, providing finer-grained and more generalizable methods for embedding heterogeneous networks. JUST (rf35) introduces the ”Jump-Stay” strategy for adaptive node sampling, further improving the representation capability. HMSG (rf36) constructs a meta-path-based multi-relational graph neural network, and employs contrastive learning together with Deep Graph Infomax (DGI) (rf37) to jointly learn high-quality node representations. HPSRW (rf38) combines higher-order Markov chains and the Spacey Walk method to effectively mitigate the stability issues in meta-path sampling. While heterogeneous graph learning has advanced in various domains, existing general-purpose models lack specialized designs for the unique characteristics of transportation networks. Even when applied to our constructed graph that explicitly incorporates OD flows and route information, these models still fail to effectively leverage its structural properties, leading to limited capability in capturing the spatial dependencies inherent in road networks. To the best of our knowledge, CRRank (rf1) is the first work to apply heterogeneous graphs to transportation network analysis, achieving node ranking by integrating OD flow and path information. However, this method relies on manually defined computational rules and lacks end-to-end learnability, leaving room for substantial improvements in model performance.
@@ 3.1 Definitions and problem statement
Definition 1 (Road Network): A road network is a directed graph $G(V^{l},{\color[rgb]{0,0,0}E^{l}})$, where $V^{l}$ is the set of nodes; $E^{l}$ is the set of directed edges; An edge $e_{i,j}$ indicates that node $v^{l}_{i}$ is the incoming neighbor of node $v^{l}_{j}$. Each node $v^{l}_{i}$ has an initial attribute vector. In this paper, $v^{l}_{i}$ specifically refers to a road segment or a grid.
@@
Definition 2 (OD Flow): An OD flow is represented as a triple $\textlangle v_{c},v_{t},f\textrangle$, where $v_{c}\in V^{l}$ denotes the origin node, $v_{t}\in V^{l}$ denotes the destination node, and $f$ represents the number of vehicles traveling from $v_{c}$ to $v_{t}$ within a specified time period. For instance, the road network in Figure 1 contains two OD flows: $v_{1},v_{4},300>$ and $v_{8},v_{7},500>$.
@@
Definition 3 (path): A path $p:\textlangle v^{l}_{1},v^{l}_{2},...,v^{l}_{|p|}\textrangle$ is a sequence of nodes in the road network, encompassing all the nodes traversed during a trip associated with a specific OD flow. Here, $|p|$ denotes the number of nodes in $p$. Evidently, an OD flow can be associated with multiple paths. For example, the OD flow $v_{8},v_{7}>$ corresponds to two paths: $v_{8},v_{9},v_{5},v_{6},v_{7}>$ and $v_{8},v_{9},v_{10},v_{11},v_{7}>$.
Definition 4 (Trip Graph): A trip graph is a heterogeneous graph $TG(V,A,M)$, where $V=V^{o}\cup V^{p}\cup V^{l}$ consists of OD nodes $(V^{o})$, path nodes $(V^{p})$, and road segment nodes $(V^{l})$ with corresponding attribute matrices $A^{o}$, $A^{p}$, and $A^{l}$ encoding features. The adjacency structure includes (1) $M_{o,p}$ representing traffic proportions between OD pairs and paths $(m_{o,p}\in[0,1])$, (2) $M_{p,l}$ as a binary path-segment incidence matrix $(m_{p,l}=\mathbb{I}(v^{l}\in p))$ and (3) a topologically weighted segment connectivity matrix $M_{l,l}$ defined as
$m_{i,j}=\left\{\begin{array}[]{c}\frac{\sum_{p\in\mathcal{P}_{i,j}}e^{-d_{i,j}^{p}}}{\sum_{p\in\mathcal{P}_{i}}\sum_{v_{k}^{l}\in p}e^{-d_{i,k}^{p}}}\text{ if }\exists p\text{ where }v_{i}^{l}\prec_{p}v_{j}^{l}\\ 0\quad\text{ otherwise }\end{array}\right$. (1)
where $\mathcal{P}_{i,j}=\left\{p\in V^{p}\mid v_{i}^{l},v_{j}^{l}\in p,v_{i}^{l}\prec_{p}v_{j}^{l}\right\}$ denotes paths containing both segments with $v_{i}^{l}$ upstream of $v_{j}^{l}$, $d_{i,j}^{p}$ counts intermediate segments between them in path $p$, $e$ is the propagation factor, and $\mathcal{P}_{i}=\left\{p\in V^{p}\mid v_{i}^{l}\in p\right\}$ is the set of paths containing $v_{i}^{l}$. Finally, the adjacency matrix $M_{p,l}$ and attribute matrices $A^{o}$, $A^{p}$, and $A^{l}$ are row-normalized using the L1 norm, resulting in $\overline{M}_{p,l}$, $\overline{A}^{o}$, $\overline{A}^{p}$, and $\overline{A}^{l}$, respectively.
@@
While the Trip Graph (TG) effectively integrates the macro-level functional organization and topological structure of the road network, it struggles to capture local, attribute-driven semantic similarities among nodes of the same type. For instance, the functional similarity between OD pairs that both connect residential and commercial areas, or between road segments sharing comparable attributes like the number of lanes or traffic capacity, is crucial for learning comprehensive and robust segment representations. To address this limitation, we introduce the Attribute-guided Graph (AG) as a complementary structure to the TG at the semantic level. Its core innovation lies in elevating node attributes to the status of explicit nodes, thereby constructing a heterogeneous bipartite graph that directly models these attribute-driven functional relationships.
@@
Definition 5 (Attribute-Guided Graph): For each node type $\phi\in\left\{o,p,l\right\}$, an attribute-guided graph is defined as a weighted bipartite graph $AG_{\phi}(V^{\phi}\cup U^{\phi},A^{\phi})$, where $V^{\phi}$ is the set of nodes of type $\phi$, $U^{\phi}$ is the set of attribute nodes associated with type $\phi$, $A^{\phi}$ is the set of weighted edges between nodes $v_{i}^{\phi}\in V^{\phi}$ and attribute nodes $u_{k}^{\phi}\in U^{\phi}$. The edge weight between $v_{i}^{\phi}$ and $u_{k}^{\phi}$ is denoted as $a_{i}^{\phi}\in A^{\phi}$. Special cases: For OD nodes $(\phi=0)$, $AG_{o}(V^{o}\cup U^{o},A^{o})$. For paths $(\phi=p)$, $AG_{p}(V^{p}\cup U^{p},A^{p})$. For road segments $(\phi=l)$, $AG_{l}(V^{l}\cup U^{l},A^{l})$.
[…]

### v3

# Learning to Rank Critical Road Segments via Heterogeneous Graphs with Origin-Destination Flow Integration

@@ Abstract
Existing learning-to-rank methods for road networks often fail to incorporate origin–destination (OD) flows and route information, limiting their ability to model long-range spatial dependencies. To address this gap, we propose HetGL2R, a heterogeneous graph learning framework for ranking road-segment importance. HetGL2R builds a tripartite graph that unifies OD flows, routes, and network topology, and further introduces attribute-guided graphs that elevate node attributes into explicit nodes to model functional similarity. A heterogeneous joint random walk algorithm (HetGWalk) jointly samples both graph types to generate context-rich node sequences. These sequences are encoded using a Transformer to learn embeddings that capture long-range structural dependencies induced by OD flows and route configurations, as well as functional associations derived from attribute similarity. Finally, a listwise ranking strategy with a KL-divergence loss evaluates and ranks segment importance. Experiments on three SUMO-generated simulated networks of different scales show that, against state-of-the-art methods, HetGL2R achieves average improvements of approximately 7.52%, 4.40% and 3.57% in ranking performance.
@@ 1 Introduction
Efficient and resilient road networks are essential for ensuring smooth urban mobility and public safety. When a single road segment becomes congested or blocked, the resulting disruption often propagates along multiple routes, leading to large-scale delays or even citywide paralysis. Therefore, identifying critical road segments—those whose failure would significantly degrade overall network performance—is of great importance for traffic management and infrastructure planning (Xu et al., 2018).
Figure 1: A small simulated road network used for illustration. The network consists of $1$ nodes ($v$) and 11 directed edges ($e$), with two origin–destination (OD) pairs: $v_{1}$, $v_{4}>$ with 300 vehicles traveling along a single path, and $v_{8}$, $v_{7}>$ with 500 vehicles distributed over two alternative paths. Red and blue arrows indicate the routing paths associated with the two OD pairs, respectively.
@@
Prior studies on road-network criticality were primarily based on complex network theory, in which centrality measures (e.g., early betweenness centrality (Freeman, 1977) and PageRank (Agryzkov et al., 2012), and more recent methods such as IDME (Sun et al., 2023), EPC (Ullah et al., 2024), OVED-Rank (Ullah et al., 2025), LSS (Ullah et al., 2023), and GLC (Ruan et al., 2025)) were used to quantify the structural importance of nodes. These approaches are intuitive and easy to interpret but fail to incorporate the rich attribute features and dynamic traffic behaviors associated with each road segment. In reality, a segment’s criticality depends on multiple factors such as traffic volume, number of lanes, and functional hierarchy, many of which are neglected in purely topological metrics. Some studies have attempted to combine centrality measures with attribute features (e.g., study (Agryzkov et al., 2012)), but most rely on manually tuned parameters, limiting their scalability and generalization to large real-world networks.
@@
In recent years, graph neural network (GNN)–based learning methods (Fan et al., 2020; Xu and Zhang, 2024) have attracted increasing attention in the field of node importance assessment. These models integrate node attributes with local topological information through message-passing mechanisms to learn effective node embeddings. However, due to the limited receptive field of GNNs, they struggle to capture the long-range path dependencies commonly present in traffic flows, which limits their performance in critical road segment ranking tasks.
@@
In transportation networks, the global influence of a road segment is largely shaped by origin–destination (OD) flows, which quantify the travel demand between each OD pair. The OD flows determine which paths are utilized and how traffic load is distributed, thereby shaping each segment’s functional role and criticality (Guo et al., 2025; Xing et al., 2025; Wang et al., 2024). For instance, if a segment lies on the shortest paths of multiple high-demand OD pairs, any incident or restriction there can induce congestion that propagates upstream along the affected OD paths, leading to severe network-wide delays. Conversely, a peripheral segment carrying a similar traffic volume but serving only a few OD flows would have a far more limited systemic impact if disrupted.
@@
Existing graph learning approaches face two fundamental limitations in this task. First, they primarily learn structural representations from the physical topology of road networks while overlooking the functional structures shaped by OD flows. Second, their local neighborhood aggregation mechanism is constrained by network depth, resulting in limited receptive fields. As a result of these limitations, they struggle to capture long-range spatial dependencies and consequently fail to accurately model the impact patterns of local disruptions (e.g., an incident at node $v_{11}$)—where the scope of impact is dictated by the functional roles of segments within OD flows and paths, rather than by physical proximity. Although CRRank (Xu et al., 2018) incorporates OD information to evaluate the functional importance of road segments, its reliance on manually designed aggregation rules and similarity metrics constrains its expressiveness and effectiveness, as it cannot learn complex interaction patterns in an end-to-end manner.
@@
    Methodological Innovation: HetGL2R constructs a tripartite graph (connecting OD pairs, paths, and road segments) to integrate the functional structure of the road network with its physical structure. Simultaneously, it introduces three attribute-node bipartite graphs—one each for OD pairs, paths, and segments—to explicitly model the attribute-based functional similarities. A joint random walk mechanism operates on these graphs collectively to sample context-rich node sequences, which are then processed by a Transformer encoder to learn the long-range dependencies and functional roles of segments for accurate criticality ranking.
@@
    Theoretical Insight: Our core insight is that spatial dependencies in traffic networks are governed by functional structure rather than physical topology. To capture this structure, we introduce a joint random walk that integrates two complementary perspectives: functional role relationships (OD-path-segment) and attribute-based functional similarities. Consequently, the model learns the underlying mechanisms of spatial dependency, thereby enabling accurate assessment of nodes’ spatial influence.
@@ 2.1 Evaluation of node importance in networks
Existing research on identifying critical nodes in networks can be broadly categorized into two approaches: centrality-based methods and machine learning-based methods. The limitation of centrality-based methods lies in their inability to fully leverage the multidimensional features of nodes. Although study (Agryzkov et al., 2012) attempts to improve critical node identification in road networks by integrating node centrality with attribute features, its performance heavily depends on manually set topology-feature trade-off parameters. This leads to a lack of generalizability in quantifying the contributions of topology and features, limiting its applicability in large-scale real-world road networks. To address these limitations, LSS (Ullah et al., 2023) estimates node importance based on local information, eliminating the need for global knowledge and complex parameter tuning. Study (Ruan et al., 2025) proposed a dual-perspective approach that identifies influential nodes by considering inter-cluster connectivity and local neighborhood importance. However, similar to traditional centrality-based methods, these approaches still fail to incorporate node attribute features or dynamic network behaviors, limiting their applicability in more complex real-world scenarios such as road networks. In recent years, researchers have gradually shifted towards machine learning-based methods, which can be divided into two main technical approaches: reinforcement learning frameworks (e.g., FINDER proposed in study (Fan et al., 2020), which utilizes GraphSAGE (Hamilton et al., 2017) to learn node importance; study (Tan et al., 2024), which combines attribute-aware graph attention networks with double DQN to optimize critical node identification in sparse graphs) and graph representation learning (e.g., TraNode2vec from study (Huang et al., 2022), which assesses node importance by integrating topological embedding clustering and traffic flow; study (Xu and Zhang, 2024), which introduces the MGL2Rank framework that employs multi-graph fusion techniques to jointly model topology and road segment attributes). However, when applying these methods to road network node importance evaluation, they fail to incorporate OD flow and path selection. As we analyzed in the introduction, these two types of information are crucial for assessing node importance in road networks. Therefore, there remains significant room for improvement in the performance of these methods.
@@ 2.2 Learning to rank
Learning to rank (LTR) originated in the field of information retrieval and is a technique for ordering documents, items, or other entities based on their relevance. In recent years, it has been widely applied in various domains, including recommender systems (Abdollahpouri et al., 2017), sentiment analysis (Song et al., 2017), and object detection (Tan et al., 2019). Depending on the characteristics of the input and output spaces, LTR methods can be broadly categorized into three types: pointwise, pairwise, and listwise approaches. The pointwise approach (Li et al., 2007) treats each item independently, formulating the ranking task as a classification or regression problem. However, this method neglects the relative relationships among items. In contrast, the pairwise approach (Burges et al., 2005) transforms the ranking task into a binary classification problem by comparing item pairs to determine their relative order. Prior research (Maurya et al., 2021) has introduced graph neural networks into pairwise ranking to approximate the importance of nodes within a network. Although the pairwise method offers improved accuracy over the pointwise method, it still fails to fully capture the global context and relative importance of items within the entire set, limiting its overall performance. The listwise approach (Xia et al., 2008b) treats ranking as a holistic optimization problem and has become a focal point of research due to its modeling and computational complexity. For example, one study (Lan et al., 2009) analyzes how list length and loss function selection affect algorithm stability through generalization error bounds, while another (Wu et al., 2016) creatively reformulates ranking as a multi-task regression problem, overcoming the limitations of traditional linear models. With the rise of deep learning, researchers (Pobrotyn et al., 2020; Ai et al., 2018; Zhu and Klabjan, 2020) have incorporated attention mechanisms and recurrent neural networks into listwise modeling, significantly enhancing performance in scenarios involving complex interactions. On the application side, DiAL (Singh et al., 2024) is the first to integrate the $\alpha$-NDCG diversity metric into listwise loss functions. Further advancements include leveraging large language models to generate reference answer sequences for listwise evaluation (Yang et al., 2024; Xiong et al., 2025), as well as developing differentiable approximations to ranking indicators that closely match true loss functions with limited data (Thonet et al., 2022). Recent work (Bai et al., 2023) also proposes a regression alignment strategy to address the multi-objective trade-off between score calibration and ranking optimization. Building upon the core principles of listwise ranking, this paper proposes a heterogeneous graph learning framework tailored for node ranking in transportation networks.
@@ 2.3 Heterogeneous graph learning
Heterogeneous graphs (Shi et al., 2018) model multiple types of nodes and edges, integrating information from heterogeneous data sources, thereby demonstrating a stronger representation ability for complex systems. Representation learning on heterogeneous graphs aims to obtain meaningful node embeddings to support various downstream tasks, such as node ranking and link prediction (Dong et al., 2012; Li et al., 2020). Many studies have focused on designing graph neural network models suitable for heterogeneous graphs. For example, most of these networks (Wang et al., 2019; Fu et al., 2020; Chang et al., 2022) utilize meta-paths and type-aware aggregation strategies to capture the heterogeneity of the graph effectively, that is, the difference in the type of nodes and edges. HetGNN (Zhang et al., 2019) combines random walk strategies with bidirectional LSTM structures, effectively integrating heterogeneous node attributes and their neighboring information. HetSANN (Hong et al., 2020) directly encodes the information of different types of nodes and edges through a type-aware attention mechanism, enhancing the model’s ability to represent the heterogeneous graph structure. These methods are distinctive in their modeling strategies, but most still rely on message-passing mechanisms, which do not fully exploit the potential structural relationships between nodes in heterogeneous graphs. To address this, M2CHGNN (Moradi et al., 2025) employs a dual-view encoding strategy that integrates attention mechanisms with contrastive learning to generate high-quality node representations for heterogeneous graphs. UI-HEHo (Xun et al., 2024) effectively learns high-order node embeddings via multi-head attention mechanisms and multi-task shared features. SchemaWalk (Samy et al., 2022) proposes an unbiased edge-type exploration strategy, providing finer-grained and more generalizable methods for embedding heterogeneous networks. JUST (Hussein et al., 2018) introduces the ”Jump-Stay” strategy for adaptive node sampling, further improving the representation capability. HMSG (Guan et al., 2023) constructs a meta-path-based multi-relational graph neural network, and employs contrastive learning together with Deep Graph Infomax (DGI) (Veličković et al., 2018) to jointly learn high-quality node representations. HPSRW (He et al., 2019) combines higher-order Markov chains and the Spacey Walk method to effectively mitigate the stability issues in meta-path sampling. While heterogeneous graph learning has advanced in various domains, existing general-purpose models lack specialized designs for the unique characteristics of transportation networks. Even when applied to our constructed graph that explicitly incorporates OD flows and route information, these models still fail to effectively leverage its structural properties, leading to limited capability in capturing the spatial dependencies inherent in road networks. To the best of our knowledge, CRRank (Xu et al., 2018) is the first work to apply heterogeneous graphs to transportation network analysis, achieving node ranking by integrating OD flow and path information. However, this method relies on manually defined computational rules and lacks end-to-end learnability, leaving room for substantial improvements in model performance.
@@ 3.1 Definitions and problem statement
Definition 1 (Road Network). A road network is modeled as a directed graph $G(V^{l},{\color[rgb]{0,0,0}E^{l}})$, where $V^{l}$ denotes the set of nodes and $E^{l}$ denotes the set of directed edges. In this paper, each node $v^{l}_{i}\in V^{l}$ represents a road segment, and a directed edge $e_{ij}^{l}\in V^{l}$ indicates a directed connection from $v^{l}_{i}$ to $v^{l}_{j}$. Each node $v^{l}_{i}$ is associated with an initial attribute vector describing its physical or traffic-related properties.
@@
Definition 2 (OD Flow). An OD flow is defined as a triple $\textlangle v_{i}^{l},v_{j}^{l},f\textrangle$, where $v_{i}^{l},v_{j}^{l}\in V^{l}$ denote the origin and destination nodes, respectively, and $f$ denotes the number of vehicles from $v_{i}^{l}$ to $v_{j}^{l}$ over a predefined observation period. For instance, the road network in Figure 1 contains two OD flows: $v_{1},v_{4},300>$ and $v_{8},v_{7},500>$.
@@
Definition 3 (Path). A path $\rho=(v^{l}_{1},v^{l}_{2},...,v^{l}_{|\rho|})$ is defined as an ordered sequence of road-segment nodes in the road network, representing the segments traversed by trips associated with a specific origin–destination (OD) flow. Here, $|\rho|$ denotes the number of road segments in path $\rho$. In practice, a single OD flow may be associated with multiple feasible paths. For example, the OD pair $v_{8},v_{7},500>$ corresponds to two paths: $(v_{8},v_{9},v_{5},v_{6},v_{7})$ and $(v_{8},v_{9},v_{10},v_{11},v_{7})$.
Definition 4 (Trip Graph). A trip graph is a heterogeneous graph $TG(V,A,M)$, where $V=V^{o}\cup V^{p}\cup V^{l}$ consists of OD nodes $V^{o}$, path nodes $V^{p}$, and road-segment nodes $V^{l}$. Each node type is associated with an attribute matrix, denoted by $A^{o}$, $A^{p}$, and $A^{l}$. The adjacency structure $M$ comprises three components: (1) OD–Path association matrix $M^{o,p}\in\mathbb{R}^{|V^{o}|\times|V^{p}|}$, where each entry $m^{o,p}\in[0,1]$ represents the traffic proportion of an OD pair assigned to a specific path; (2) Path–Segment incidence matrix $M^{p,l}\in\left\{0,1\right\}^{|V^{p}|\times|V^{l}|}$, where $m^{p,l}=\mathbb{I}(v^{l}\in\rho)$ indicates whether road segment $v^{l}$ is contained in path $\rho$. (3) Segment–Segment connectivity matrix $M^{l,l}\in\mathbb{R}^{|V^{l}|\times|V^{l}|}$, whose elements are defined as
$m_{i,j}^{l,l}=\sum_{\begin{subarray}{c}\rho\in(\mathcal{P}_{i}\cap\mathcal{P}_{j})\\[2.0pt] v_{i}^{l}\prec_{\rho}v_{j}^{l}\end{subarray}}e^{-c_{i,j}^{\rho}}$ (1)
where $\mathcal{P}_{i}$ and $\mathcal{P}_{j}$ denote the sets of paths containing $v_{i}^{l}$ and $v_{j}^{l}$, respectively. The notation $v_{i}^{l}\prec_{\rho}v_{j}^{l}$ indicates that segment $v_{i}^{l}$ lies upstream of $v_{j}^{l}$ along path $\rho$, $c_{i,j}^{\rho}$ is the number of intermediate segments between them on $\rho$. If no such path exists, $m_{i,j}^{l,l}=0$. $e$ is a propagation decay factor. This formulation models the strength of influence propagation between segments along OD-induced paths.
@@
While the Trip Graph (TG) effectively integrates the macro-level functional organization and physical structure of the road network, it is less effective at capturing local, attribute-based semantic similarities among nodes of the same type. For instance, the functional similarity between OD pairs that both connect residential and commercial areas, or between road segments sharing comparable attributes like the number of lanes or traffic capacity, is crucial for learning comprehensive and robust segment representations. To address this limitation, we introduce the Attribute-Guided Graph (AG) as a complementary semantic structure to the TG. Its core innovation lies in explicitly representing node attributes as nodes, thereby constructing a heterogeneous bipartite graph that directly models these attribute-based functional relationships.
@@
Definition 5 (Attribute-Guided Graph). For each node type $\phi\in\left\{o,p,l\right\}$, where $o$, $p$ and $l$ denote OD pairs, paths, and road segments respectively, an Attribute-Guided Graph is defined as a weighted bipartite graph $AG_{\phi}(V^{\phi}\cup U^{\phi},A^{\phi})$. Here, $V^{\phi}$ denotes the set of nodes of type $\phi$, $U^{\phi}$ denotes the set of attribute nodes associated with type $\phi$, and $A^{\phi}$ denotes the set of edges connecting $v_{i}^{\phi}\in V^{\phi}$ and $u_{k}^{\phi}\in U^{\phi}$. Each edge $(v_{i}^{\phi},u_{k}^{\phi})\in A^{\phi}$ is assigned a weight $a_{ik}^{\phi}$, which reflects the similarity between node $v_{i}^{\phi}$ and attribute $u_{k}^{\phi}$. Special cases include $AG_{o}(V^{o}\cup U^{o},A^{o})$ for OD pairs, $AG_{p}(V^{p}\cup U^{p},A^{p})$ for paths, and $AG_{l}(V^{l}\cup U^{l},A^{l})$ for road segments.
Definition 6 (Learning to Rank Road Segments). Let each road segment $v_{i}^{l}\in V^{l}$ be represented by a feature vector $h_{i}$. Given a list of $K$ road segments with representations $H=\left[h_{1},h_{2},\cdots,h_{K}\right]$, and their corresponding ground-truth importance ordering $Y=\left[y_{1},y_{2},\cdots,y_{K}\right]$, where $y_{i}$ denotes the rank position of segment $v_{i}^{l}$ and higher-ranked segments are considered more critical, the learning-to-rank task aims to learn a parametric mapping $F_{\theta}:\mathbb{R}^{K\times d}\rightarrow Y^{K}$, where $Y^{K}$ denotes the set of all possible permutations of $K$ elements. The output $\hat{Y}=F_{\theta}(H)$ represents the predicted ranking of the input road segments. The parameters $\theta$ are learned by minimizing a listwise ranking loss that measures the discrepancy between the predicted ranking $\hat{Y}$ and the ground-truth ordering $Y$.